Sections of the textbook covered

• All of Chapter 1, except the Hill Cipher.

• All of Chapter 2, except Huffman encodings and product cryptosystems.

• Chapter 3.

• All of Chapter 4, except the p − 1 method for factoring.

• Sections 5.1, 5.2.2, 5.3, and 5.4.

• Sections 6.1, 6.2, 6.3, and 6.5.

• Section 7.1, 7.2, and 7.3.

• Section 8.4.

• Section 11.1.

• Section 12.4.

• Chapter 13.